5. 6. 7. 8. 9.

5.

media%2F178%2F178f6458-44db-4e72-b47a-ac

6.

media%2Faad%2Faad726a2-096e-41f4-9e9b-a6


media%2Fdb0%2Fdb0d8cea-a748-4cfe-9a21-fd

7.

media%2F544%2F5440e57e-6ec0-4dc8-b6b0-5f


8.

media%2F63f%2F63f91769-9781-4119-a6b4-14


9.

media%2F493%2F493db679-aa3e-470f-9b99-0e


media%2Fe43%2Fe436ab2c-f37a-427f-8a7f-72
Answer #1

Multiple questions posted and i am solving first one
5)for 2018 the journal entries are
Income tax expense(db)64
Deferred tax asset(db)100*40%=40
Income tax payable(cr)260*40%=104mn
The deferre potion in 2019 is 90 and in 2018 it is 100 so the difference is 10mn

for 2019 the entries are:
Income tax expense(db)108mn
Deferred tax asset(db)10*40%=4mn
Income tax payable(cr)280*40%=112mn